The total concentration of dissolved ions in seawater is about 1.1 M, but concentrations can vary depending on the exact location and conditions. What minimum pressure must be supplied to a reverse osmosis water purification system in order to overcome the osmotic pressure in water that is 1.06 M in total dissolved ions at 34 C?
1
Expert's answer
2019-09-16T04:02:55-0400

The osmotic pressure in seawater with 1.06 M of total ions concentration would be calculated using vant Hoff expression:


"p=iCRT,"


where i is equal to "1-a(n-1)".

As for the seawater i or vant Hoff factor is nearly equal to the number of ions produced by one formula unit of salt, like NaCl etc. Thus, i "\geq" 2 (NaCl = Na+ + Cl-, two ions).

Now it is the right time to calculate the minimum pressure required:

"p \geq 2*1060 \frac{mol}{m^3}*8.31 \frac{J}{mol*K}307K=2.7MPa".